Truball Inc., which manufactures sports equipment, consists of several operating divisions. Division A has decided to go outside the company to buy materials since division B plans to increase its selling price for the same materials to $200. Information for division A and division B follows:

Outside price for materials $185
Division A’s annual purchases 13,500 units
Division B’s variable costs per unit $175
Division B’s fixed costs, per year $ 1,320,000
Division B’s capacity utilization 100

%

Required:

1. Assume that division B cannot sell its materials to outside buyers. Calculate the net cost or benefit to the company as a whole if Division A purchases the materials outside the company.

2-a. Assume that division B can save $245,000 in fixed costs if it does not manufacture the material for Division A. Calculate the net cost or benefit to the company as a whole for A to purchase outside the company.

2-b. From the standpoint of the effect of the transaction on the company as a whole, should Division A purchase from the outside market?

3-a. Assume the situation in Requirement 1. If the outside market value for the materials drops $12, calculate the net cost or benefit to the company as a whole for A to purchase outside the company.

3-b. From the standpoint of the effect of the transaction on the company as a whole, should Division A purchase from the outside market?

I have 1, 2B, 3A but getting incorrect answers for 2A and 3A, so I need help on those. Please help!
Answer #1

Since Division B cannot sell outside, cost to the company = (185-175)*13500

= $135,000

Since fixed costs will not be avoided

2-a. Benefit of purchasing outside = Savings in fixed costs – Price paid over variable cost

= 245,000-135,000

= $110,000

2-b. Yes, should purchase since benefit

3-a.Benefit = (175-173)*13,500 = $27000

3-b. Yes, since benefit